OpenStudy (anonymous):

Is this right? Q: Solve 13+s/p=2 for p A: p(13)=2(13)+2=28? What is s? Ill say 2. 28/13≈2.15

OpenStudy (unklerhaukus):

nope

OpenStudy (anonymous):

Which part isnt right?

OpenStudy (anonymous):

Goddammit this is just hopeless..

OpenStudy (turingtest):

hi Inopeki do not despair! practice practice...

OpenStudy (turingtest):

what made you feel like putting a 13 there? we were supposed to get a function like before

OpenStudy (anonymous):

And didnt i?

OpenStudy (turingtest):

you can't start plotting points like that until we've solved it for one of the variables. The answer to the question I was looking for should have 2 variables.

OpenStudy (turingtest):

13+s/p=2 what do you get multiplying both sides by p? do nothing else but that.

OpenStudy (anonymous):

I can go check

OpenStudy (turingtest):

no, I was wrong, nevermind

OpenStudy (unklerhaukus):

13+s/p=2 for p 11+s/p=0 11=-s/p p=-s/11

OpenStudy (anonymous):

See? This guy knows it. Why cant i?

OpenStudy (anonymous):

Whatever. give me a new one

OpenStudy (unklerhaukus):

You can with Time Inopeki

OpenStudy (turingtest):

well, there's one way. btw Rhauckus I made up this problem for Inopeki to teach him how to solve for a variable @Inopeki, he has many years doing this, you need to practice basics. It will be like that for a while.

OpenStudy (anonymous):

Well that sucks.. Still, if i really need to learn this i guess i will. Give me a new example so i can practise please

OpenStudy (turingtest):

a new one? solve\[A=2\pi r h\] for A, r, and h I need to get you to stop thinking in numbers.

OpenStudy (turingtest):

that's three problems, by the way.

OpenStudy (anonymous):

is \[\pi\] 3.14 here or is it a variable?

OpenStudy (turingtest):

no, it's about 3.14159..., that's why I didn't make it something to solve for. but don't think of it as a number, leave it as pi. The number does not matter.

OpenStudy (turingtest):

A=2(pi)rh

OpenStudy (anonymous):

Ok, so the answer should not be a number?

OpenStudy (turingtest):

not at all

OpenStudy (anonymous):

Interesting..

OpenStudy (turingtest):

it should be a function like r=5t+4 it has some numbers, but not all.

OpenStudy (anonymous):

One question. If i, in the answer have to give you As "definition", how can that not be a number?

OpenStudy (turingtest):

The formula I gave you is for the surface area of the side of a cylinder. The surface area of any particular cylinder depends on its radius and length. So the formula covers all possible cylinders, but only for some particular cylinders will you get a certain value for the surface area. It is a function, this time of both r and h.

OpenStudy (turingtest):

if I tell you the radius and height of our particular cylinder, only then can you get a number.

OpenStudy (anonymous):

I just studied that in school! I thought we were talking algebra here..

OpenStudy (akshay_budhkar):

@Turning you train guys well I will need your help for multivariable calculus, Laplace and stuff I will let you know then :D Great job is all I can say :D

OpenStudy (turingtest):

Solving for a variable is algebra. They are all connected as you will learn. Thanks Akshay!

OpenStudy (anonymous):

But that still doesnt answer my question, how should i define A?

OpenStudy (turingtest):

in this case, A is already defined by\[A=2\pi rh\]that is basically the definition of the surface area of the side of a circular cylinder. No more definition is needed. I wanted you to catch that the first problem, solve for A, is already done for you: we already have A=(something) so it IS solved for A. So now solve it for r, if you can. what can we divide by to get r by itself?

OpenStudy (anonymous):

We can divide h with A, right?

OpenStudy (anonymous):

No.. that cant be it.

OpenStudy (turingtest):

when you say 'divide h with A' I'm not sure what you mean. Do you mean divide both sides by A or divide both sides by h ??? One is okay, one is not. You know that division is the inverse of multiplication. So ask yourself 'what is r being multiplied by?' and divide by that. Well, what is r being multiplied by?

OpenStudy (anonymous):

The height.

OpenStudy (turingtest):

is that all?

OpenStudy (anonymous):

Actually no

OpenStudy (turingtest):

what else?

OpenStudy (anonymous):

pi^2

OpenStudy (anonymous):

right?

OpenStudy (turingtest):

where did you see a squared term?

OpenStudy (anonymous):

I have no idea.. i realize thats wrong now cause that is for getting the volume of a cylinder..

OpenStudy (anonymous):

And that concludes that i am an utter and complete retard

OpenStudy (turingtest):

it's okay. so again, what, in all, is r being multiplied by?

OpenStudy (turingtest):

c'mon, being hard on yourself never helped anything.

OpenStudy (anonymous):

2 and then pi

OpenStudy (turingtest):

and...?

OpenStudy (anonymous):

height?

OpenStudy (anonymous):

wait!

OpenStudy (turingtest):

right, it's being multiplied by 2(pi)h so that is what we need to divide by no don;t take it back!

OpenStudy (anonymous):

if you divide both sides by h, and both sides by pi you should isolate it so you have 2r, right?

OpenStudy (turingtest):

true!

OpenStudy (anonymous):

Yeah.. pretty much what you said

OpenStudy (anonymous):

But that doesnt mathematically give us anything..

OpenStudy (anonymous):

Then you just divide both sides by 2 and get r

OpenStudy (turingtest):

sure it does, it gives us the answer not all answers in math are numbers, in fact I hate using numbers if I can help it. about your last post, yes, so what do we get for r=?

OpenStudy (anonymous):

What do you mean?

OpenStudy (turingtest):

what is on the other side of the equation when we solve for r?

OpenStudy (turingtest):

what did we divide both sides by?

OpenStudy (anonymous):

2(åi)h?

OpenStudy (anonymous):

(pi)*

OpenStudy (turingtest):

right. so what will A be divided by?

OpenStudy (anonymous):

well we get r on both sides right?

OpenStudy (anonymous):

2(pi)h

OpenStudy (turingtest):

no, just one and all alone, that's what it means to solve for r again, yes to your last post so write out r=...?

OpenStudy (anonymous):

1

OpenStudy (turingtest):

where did you get 1 ?

OpenStudy (anonymous):

Nevermind, my mind is messing up my mind, if you know what i mean

OpenStudy (turingtest):

If you are tired then take a break here is the answer, watch is closely: \[A=2\pi rh\]divide both sides by\[2\pi h\]giving\[r=\frac{A}{2\pi h}\]If you want to keep going then try just something simpler with NO numbers:\[ab=c\]solve for a

OpenStudy (anonymous):

Maybe i should catch some sleep, its 4:50 am here..

OpenStudy (turingtest):

that is a very good idea then. sleep on it and try it in the morning. I can't tell you how many problems I've solved just by sleeping on them. Don't worry, step-by-step. You will get there I know. Goodnight!

OpenStudy (anonymous):

Good night, thank you for all the help!

OpenStudy (turingtest):

anytime :)

OpenStudy (anonymous):

A=2πrh divide both sides by 2πh giving r=A2πh If you want to keep going then try just something simpler with NO numbers: ab=c solve for a When you say "solve for a", do you mean i should isolate it? I guess we would divide both sides by b and then c.

OpenStudy (anonymous):

Actually, just b.

OpenStudy (anonymous):

I would then get a/b=c/b Simplified, a=c/b

OpenStudy (turingtest):

^^^correct! now get some sleep dude! the mind needs to be fresh.

OpenStudy (anonymous):

I did sleep, that answer was today.

OpenStudy (turingtest):

lol I thought you stayed up all night. Told you sleeping on it would help. now solve ab=c for b

OpenStudy (anonymous):

Thats easy, both sides divided by a, you get b/a=c/a Simplify: b=c/a

OpenStudy (anonymous):

no!

OpenStudy (turingtest):

yes^^

OpenStudy (anonymous):

actually, yes

OpenStudy (turingtest):

that's what I wanted to hear since it's so easy A=2πrh for r

OpenStudy (anonymous):

Divide both sides with 2(pi)h, you get A/2(pi)h=2(pi)rh/2(pi)h Simplified: A/2(pi)h=r

OpenStudy (turingtest):

awesome! I know you got it but you should write it as A/(2π)h=r parentheses make it clear what is on the bottom.

OpenStudy (anonymous):

Yes! Oh ok.

OpenStudy (turingtest):

sorry, meant A/(2πh)=r

OpenStudy (turingtest):

anyway, slightly different one: solve 5t=t/2+s for s, then for t

OpenStudy (anonymous):

So the whole bottom part in parenthesis?

OpenStudy (turingtest):

that's two problems^^^

OpenStudy (anonymous):

Multiply both sides by 2+s, right?

OpenStudy (turingtest):

right, because everything in parentheses is on the bottom if you wrote 1/2x I don't know whether you mean\[\frac{1}{2x}\]or\[\frac{1}{2}x\]so parentheses are nice to clear things up think you can do the other? solve \[5t=\frac{t}{2}+s\] for s a good first step is to subtract t/2 from both sides. Then s will be isolated, right?

OpenStudy (anonymous):

Oh, i thought you meant t/(2+s)

OpenStudy (turingtest):

I would have used parentheses if I meant that. if I had meant that then you would have been right, so you'll get one like that soon. but for now solve it as it is for s

OpenStudy (anonymous):

Right, so now i have 5t-t/2=s, am i right?